carbon tetrachloride CCl_4 occupies a volume of 0.001 m^3 with a mass of 1.6 kg at a temp of 20^o C . Determine its density and specify volume at given temp?

1
Expert's answer
2022-05-23T16:24:35-0400

"\rho=\frac mV=1600~\frac{kg}{m^3},"

"v=\frac 1\rho=625~\frac{cm^3}{kg}."